# A Hypernetwork-Based Approach to KAN Representation of Audio Signals

### v1

@@ Results
2 23,016 0.0033 30.7170 2.3927 9.0285 1.4980 0.7556 0.3731
4 25,704 0.0017 33.7562 1.6014 12.3093 1.9463 0.9040 0.2561
8 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
10 33,768 0.0005 41.1152 1.3921 20.0078 3.4947 0.9930 0.1574
12 36,456 0.0005 42.6504 1.4182 21.6062 3.5176 0.9949 0.1664
16 41,832 0.0002 46.8576 1.4194 25.9786 3.6256 0.9957 0.1538
24 52,584 0.0003 45.7863 1.4764 24.9128 3.6033 0.9955 0.1982
Table 7: KAN results on 1.5 s samples in relation to different number and sizes of hidden layers.
@@
[24, 12, 6] 10,500 0.0014 35.9192 1.4806 14.5340 2.1313 0.9360 0.2977
[24, 24, 12, 6] 18,564 0.0012 37.6839 1.3589 16.4163 2.7035 0.9778 0.2746
[48, 12, 6] 19,908 0.0010 38.5736 1.3589 17.3090 2.8569 0.9795 0.2679
[48, 12, 12, 6] 21,924 0.0010 38.2946 1.3557 17.0775 2.9597 0.9810 0.2618
[48, 16, 8] 23,408 0.0009 38.7324 1.3347 17.5094 3.0476 0.9845 0.2451
[48, 24, 12] 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
[96, 48, 24] 102,480 0.0003 43.6428 1.1715 22.5679 4.0203 0.9964 0.1072
We further analyze the influence of various parameters’ values on the KAN’s performance. In the ablation study, experiments are also conducted on audio samples of 1.5 seconds and 3 seconds, with each subset limited to only 10 recordings. Due to the page limit, we include selected results for 1.5 s audio. More parameters and analysis of 3 s audio samples can be found in Appendix A. During our study we keep the following base parameters’ values, changing only the analyzed components: encoding length = 8, grid size = 10, spline order = 2, layer sizes = [48, 24, 12], scale noise = 0.1, scale base = 1, scale spline = 1, learning rate = 5e-3 and enable scale spline = True, which controls whether we include the learnable spline weigths matrix $w_{s}$. Theoretically, these weigths are redundant, as they could be absorbed into the spline itself.
Table 6 illustrates the impact of the length of the encoding (size of the positional encoding) that directly influences the size of networks. Values smaller than the baseline of 8 result in progressively worse performance; meanwhile, increasing the length provides improvement up to value of 16, where the results are the best for all but one metric (LSD).
Table 7 demonstrates the impact of different hidden layer configurations. The results indicate a consistent performance improvement as network capacity increases, with the architecture containing hidden layers of [96, 48, 24] achieving the best performance. Decreasing the networks size leads to progressive results degradation.
Table 8 presents a comparison of the varying spline order values (the degree of the applied spline). Increasing the spline order allows creation of more complex spline functions; however, this comes at the cost of additional parameters. In our study, an order of 3 provides the best overall performance across both quantitative (PSNR) and perceptual (PESQ) metrics. Using smaller or larger spline orders leads to a decline in metric scores.
Table 9 provides insights into the influence of grid size (number of knots defining the spline) on KAN’s learning process. Similarly to encoding length and spline order, an increase in grid size leads to a higher number of INR parameters. While a larger spline order eventually results in a decline in reconstruction quality, this effect is not observed with grid size. Extra parameters enhance the expressiveness of the INR, improving all metrics except LSD.
Table 8: KAN spline order (O) for 1.5 s audio recordings
@@
Table 9: Impact of spline function grid size (G) on 1.5 s reconstructions
@@ G Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
The spectrograms shown in Fig. 2 illustrate differences in the fidelity of sound reconstruction achieved by various INRs compared to the ground truth. Across all methods, the harmonic structures in the low-frequency range are generally well-preserved, reflecting their ability to capture fundamental features of the audio signal. However, significant variations emerge in the representation of high-frequency components, where most INRs lowers the intensity of the components. It is particularly visible with FINER. Moreover, all of the INRs (with KAN and NERF to the least degree) demonstrate a tendency to over-smooth the higher intensities in high-frequency components, which results in adding white noise. Finally, one can note the lower average intensity in the reconstructed audio.
@@
Figure 2: The comparison of spectrograms. We compare the original sample with the waveforms reconstructed using different INR approaches.
5 Conclusion
@@
This work demonstrates the effectiveness of Kolmogorov-Arnold Networks (KANs) as implicit neural representations for audio signals, offering high reconstruction accuracy when integrated with hypernetwork-based architectures. By evaluating KANs against classical INR models like NeRF, FINER, WIRE, RFF, and SIREN, we highlight their potential as a foundational approach for efficient audio modeling, resulting in top-2 mean in 5 out of 7 considered quality metrics for 1.5 s audio, and top-2 in all 7 for 3 s audio. Additionally, by integrating KAN into hypernetwork frameworks like FewSound, performance improves significantly, surpassing HyperSound with a 33.3% reduction in MSE, a 60.87% increase in SI-SNR, and an 8.66% in PESQ. The results suggest that KANs can serve as a powerful tool for audio representation, enabling adaptive processing across various datasets and tasks.
@@
Since our experiments were conducted at a fixed sampling rate of 22,050 Hz, the impact of varying sampling strategies requires investigation in future work. Furthermore, integration with modern neural audio codecs and speech encoder systems presented in Tab 5 might be further expanded, in particular when investigating multilingual approach. The evaluation was restricted to English datasets due their prevalence. Additionally, the datasets lacked complex auditory scenarios, such as multi-speaker dialogues, emotional speech, and diverse acoustic scenes, which might be an interesting extensions to the research. It is important to note, that while KANs demonstrated efficiency in certain tasks, their computational demands may grow significantly for longer sequences or large-scale applications.
@@
Future research should extend KAN-based models to multilingual datasets and diverse linguistic settings to enhance their generalization. Investigating the impact of sample length on performance could provide deeper insights into the balance between computational efficiency and fidelity. Architectural refinements may reduce resource requirements while maintaining accuracy, making KANs more scalable for practical deployment. Exploring their integration with multimodal frameworks, such as facial expressions or gestures, could enable advanced audio-visual synthesis and applications like avatar generation, an interesting path given hypernetworks already being used in visual modality (Przewiezlikowski et al., 2024; Kania et al., 2024). Additionally, developing interpretable models would improve transparency and understanding of KANs’ decision-making processes, which would increase trust in real-world applications and meeting legal requirements for practical usage. Finally, optimizing compatibility with a broader range of audio encoding systems could significantly enhance reconstruction quality, expanding the practical utility of KAN-based audio representations.
Appendix A Extended ablation study
At the beginning of this section we conclude the ablation study for 1.5 s audio samples by analyzing the influence of learning rate, standalone scale spline flag, scale base factor, scale noise factor and scale spline factor on KAN’s performance.
Table 10: Effect of various learning rate values on 1.5 s recordings
Learning rate Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@ Appendix A Extended ablation study
0.01 31,080 0.0007 40.5596 1.3698 19.4728 3.2799 0.9877 0.1919
0.007 31,080 0.0007 40.5170 1.3284 19.3607 3.3883 0.9908 0.1927
0.006 31,080 0.0007 40.0159 1.3051 18.9209 3.3532 0.9911 0.1770
0.005 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
0.001 31,080 0.0011 37.8920 1.3626 16.6339 2.5975 0.9711 0.2314
Table 10 shows the comparison of learning rates. Generally, KAN performs well and stably with higher learning rate setting. A learning rate of 0.001 yields the worst scores across all entries in the table. A learning rate of 0.01 achieves the best results in the PSNR and Si-SNR metrics, which are quantitative, while a learning rate near 0.006 produces the highest perceptual scores.
Table 11 shows the comparison of performance of KAN depending on whether the enable scale spline flag is set to False or True. The former results in skipping the theoretically redundant weigths. However, it leads to a drop in all metric scores except MSE, with perceptual metrics being affected the most.
Table 11: Enable scale spline flag ablation for 1.5 s audio source
Enable Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
False 28,860 0.0007 39.4067 1.5253 18.2321 2.7388 0.9653 0.2771
True 31,080 0.0007 40.0990 1.2884 18.9503 3.3588 0.9904 0.1913
At the end of this paragraph, we consider hyperparameters related to KAN initialization. The weight matrices $w_{b}$ and $w_{s}$ are initialized according to Kaiming uniform schema - the vaules are drawn from $\mathcal{U}(-k,k)$ distrubution, where
$k=\sqrt{\frac{6}{(1+5a^{2})\cdot fan\_in}}$
and $a$ is equal to scale base or scale spline for $w_{b}$ and $w_{s}$ respectively. The $c_{i}$s parameters are initialized by sampling a noise from $\mathcal{N}(0,\sigma^{2})$ distribution, and solving an instance of the least squares problem, so that $spline(x)$ for a certain $x$ is as close to the sampled noise as possible. The scale noise parameter is the standard deviation $\sigma$ of the noise distribution.
Tables 12 and 13 show the ablation results for scale spline and scale base parameters respectively. In case of the first one, 0.5 seems to be the most optimal choice with the highest score for all metrics excluding LSD, where optimal choice is 1, and STOI, where the difference between the best scores is negligible. In case of scale spline the choice of the best value varies between metrics. Table 14 shows the results for the study of scale noise hyperparameter, with the lowest tested value of 0.025 attaining the best results in most of the metrics.
Table 12: KAN performance on 1.5 s, with respect to the scale base values.
Scale Base Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
0.25 31,080 0.0008 39.9112 1.3386 18.7366 3.3149 0.9888 0.1872
0.5 31,080 0.0006 41.0193 1.3111 19.8693 3.4322 0.9904 0.1875
1 31,080 0.0006 40.8162 1.2895 19.6861 3.3821 0.9900 0.1930
2 31,080 0.0006 40.5262 1.3129 19.4204 3.2141 0.9906 0.2013
4 31,080 0.0008 39.5773 1.3195 18.4513 3.2608 0.9876 0.2062
Table 13: KAN performance on 1.5 s, with respect to the scale spline values.
Scale Spline Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
0.25 31,080 0.0006 40.6718 1.3196 19.5770 3.5030 0.9917 0.1932
0.5 31,080 0.0005 40.9010 1.3057 19.8454 3.4086 0.9915 0.1895
1 31,080 0.0007 40.5384 1.2895 19.4369 3.4281 0.9915 0.1872
2 31,080 0.0006 41.5030 1.2973 20.3684 3.3479 0.9914 0.1815
4 31,080 0.0006 40.8259 1.2981 19.7534 3.3932 0.9903 0.1887
Table 14: KAN performance on 1.5 s, with respect to the scale noise values.
KAN Scale Noise Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
0.4 31,080 0.0007 40.9740 1.3010 19.8608 3.4433 0.9910 0.1896
0.2 31,080 0.0007 40.2776 1.2849 19.1855 3.4280 0.9915 0.1796
0.1 31,080 0.0007 40.5541 1.2796 19.4436 3.4576 0.9913 0.1897
0.05 31,080 0.0007 40.0844 1.2864 18.9432 3.4004 0.9909 0.1846
0.025 31,080 0.0006 41.0045 1.2878 19.9189 3.4605 0.9897 0.1825
Performance analysis of KAN on 3 s sounds
Next, we present the ablation study for 3 s recordings, considering the same hyperparameters as used for the 1.5 s samples.
Table 15 shows the impact of different encoding lengths on 3 s utterances. The absolute performance is generally lower across all metrics in relation to 1.5 s samples, e.g. the best PESQ decreasing from 3.62 to 2.90 and the best SI-SNR decreasing from 25.97 to 17.17. The results are much less conclusive than in case of 1.5 s, where length 16 was clearly optimal across most metrics. However, the majority of metrics (MSE, PSNR, SI-SNR and PESQ, STOI) are the best for encoding lengths of 24 and 16 respectively.
Table 15: KAN performance on the 3 s samples, with respect to the encoding length.
Encoding Length Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
2 23,016 0.0036 30.4569 3.3098 4.2598 1.2121 0.6627 0.4217
4 25,704 0.0011 35.4080 2.0309 10.3129 1.5506 0.8204 0.2926
8 31,080 0.0006 38.9270 1.2706 14.0260 2.4971 0.9689 0.3578
10 33,768 0.0006 39.3990 1.3152 14.5468 2.8272 0.9779 0.3795
12 36,456 0.0005 39.6602 1.3722 14.7926 2.7818 0.9799 0.3730
16 41,832 0.0004 40.7796 1.4383 15.9502 2.9095 0.9820 0.3516
24 52,584 0.0003 41.9264 1.6033 17.1797 2.8385 0.9816 0.3429
Table 16 shows how sizes and depths of hidden layers influence results on 3 s samples. Similarly to 1.5 s (Table 7) the best performance is achieved by the largest network composed of [96, 48, 24] layers. The performance, however does not improve with the increasing size as gradually as for 1.5 s samples.
Table 16: KAN results on 3 s samples in relation to different number and sizes of hidden layers.
Layer Sizes Num Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM

### v2

@@ 4.3. KAN Ablation Study
2 23,016 3.3e-3 30.72 2.39 9.03 1.50 0.76 0.37
4 25,704 1.7e-3 33.76 1.60 12.31 1.95 0.90 0.26
8 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
10 33,768 5e-4 41.12 1.39 20.01 3.49 0.99 0.16
12 36,456 5e-4 42.65 1.42 21.61 3.52 0.99 0.17
16 41,832 2e-4 46.86 1.42 25.98 3.63 1.00 0.15
24 52,584 3e-4 45.79 1.48 24.91 3.60 1.00 0.20
Table 8. KAN results on 1.5 s samples in relation to different number and sizes of hidden layers.
@@
[24, 12, 6] 10,500 1.4e-3 35.92 1.48 14.53 2.13 0.94 0.30
[24, 24, 12, 6] 18,564 1.2e-3 37.68 1.36 16.42 2.70 0.98 0.27
[48, 12, 6] 19,908 1e-3 38.57 1.36 17.31 2.86 0.98 0.27
[48, 12, 12, 6] 21,924 1e-3 38.29 1.36 17.08 2.96 0.98 0.26
[48, 16, 8] 23,408 9e-4 38.73 1.33 17.51 3.05 0.98 0.25
[48, 24, 12] 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
[96, 48, 24] 102,480 3e-4 43.64 1.17 22.57 4.02 1.00 0.11
Next, we analyze the influence of various parameters’ values on the KAN’s performance. In this ablation study, experiments are also conducted on audio samples of 1.5 seconds and 3 seconds, with each subset limited to only 10 recordings. During our study we keep the following base parameters’ values, changing only the analyzed components: encoding length = 8, grid size = 10, spline order = 2, layer sizes = [48, 24, 12], learning rate = 5e-3 and enable scale spline = True, which controls whether we include the learnable spline weigths matrix $w_{s}$. Theoretically, these weigths are redundant, as they could be absorbed into the spline itself. In this section, we focus on the 1.5s samples and the most important parameters. The remaining ablation studies can be found in section A.1.
Table 7 illustrates the impact of the length of the encoding (size of the positional encoding), one of the parameters directly influencing the size of the network. Values smaller than the baseline of 8 result in progressively worse performance; meanwhile, increasing the length provides improvement up to value of 16, where the results are the best for all but one metric (LSD).
Table 8 shows that increasing hidden layer capacity consistently improves performance, with the [96, 48, 24] configuration performing best. Smaller architectures yield progressively worse results.
Furthermore, Table 9 presents a comparison of the varying spline order values (the degree of the applied spline). Increasing the spline order allows creation of more complex spline functions; however, this comes at the cost of additional parameters. In our study, an order of 3 provides the best overall performance across both quantitative (PSNR) and perceptual (PESQ) metrics. Using smaller or larger spline orders leads to a decline in metric scores.
Finally, Table 10 provides insights into the influence of grid size (number of knots defining the spline) on KAN’s learning process. Similarly to encoding length and spline order, an increase in grid size leads to a higher number of INR parameters. While a larger spline order eventually results in a decline in reconstruction quality, this effect is not observed with grid size. Extra parameters enhance the expressiveness of the INR, improving all metrics except LSD.
Table 9. KAN spline order (O) for 1.5 s audio recordings
@@
Table 10. Impact of spline function grid size (G) on 1.5 s reconstructions
@@ 4.4. Frequency-based analysis
The spectrograms shown in Fig. 3 illustrate differences in the fidelity of sound reconstruction achieved by various INRs compared to the ground truth. Across all methods, the harmonic structures in the low-frequency range are generally well-preserved, reflecting their ability to capture fundamental features of the audio signal. However, significant variations emerge in the representation of high-frequency components, where most INRs lower the intensity of the components. It is particularly visible with FINER. Moreover, all of the INRs (with KAN and NERF to the least degree) demonstrate a tendency to over-smooth the higher intensities in high-frequency components, which results in adding white noise. Finally, one can note the lower average intensity in the reconstructed audio.
@@
Figure 3. The comparison of spectrograms. We compare the original sample (GT) with the waveforms reconstructed using different INR approaches.
Spectral Analysis
To demonstrate KAN’s improvements over previous INRs in high-frequency information of the reconstruction, we designed an additional experiment. We can treat the absolute values of the FFT of the audio signal as probability distribution of the base frequencies. Then, we can compare distributions of the ground truth and the reconstruction using Wasserstein distance. By using $[0,1,4,...(len(audio)-1)^{2}]$ as the support of the sample instead of intuitive $[0,1,2,...len(audio)-1]$, we can make the metric more biased towards high frequencies. Using this metric in our standard experimental setting for INRs, we obtain the following results, presented in Table 11. They indicate that KAN achieves the lowest score of 79.72 among all baseline methods. The second-best outcome is obtained by WIRE, while RFF performs the worst.
Table 11. Wasserstein Distance (WD) of reconstruction and ground truth frequency distributions with the high-frequency biased support. For readability purposes the actual values are scaled down by a constant factor of 32768, that is the length of the sample (lower score is better).
NERF SIREN FINER WIRE RFF KAN
WD 184.79 218.54 205.75 131.56 362.98 79.72
5. Conclusion
@@
This work demonstrates the effectiveness of Kolmogorov-Arnold Networks (KANs) as implicit neural representations for audio signals, offering high reconstruction accuracy when integrated with hypernetwork-based architectures. By evaluating KANs against classical INR models like NERF, FINER, WIRE, RFF, and SIREN, we highlight their potential as a foundational approach for efficient audio modeling, resulting in top-2 mean in 5 out of 7 considered quality metrics for 1.5 s audio, and top-2 in all 7 for 3 s audio. Additionally, we introduce a new hypernetwork framework called FewSound, and show that by integrating KAN performance improves significantly, surpassing HyperSound with a 33.3% reduction in MSE, a 60.87% increase in SI-SNR, and an 8.66% in PESQ. The results suggest that KANs can serve as a powerful tool for audio representation, enabling adaptive processing across various datasets and tasks.
@@
As the experiments used a fixed 22,050 Hz sampling rate, exploring alternative sampling strategies remains a future work. Integration with modern neural codecs and speech encoders (Tab. 5) could be extended, particularly for multilingual scenarios. Moreover, our evaluation focused on English due to dataset availability and lacked complex auditory conditions (e.g., multi-speaker, emotional speech, diverse scenes), suggesting avenues for future research. It is important to note that while KANs demonstrated efficiency in certain tasks, their computational demands may grow significantly for longer sequences or large-scale applications.
@@
Future research should extend KAN-based models to multilingual datasets and varied linguistic contexts to improve generalization. Analyzing the effect of sample length may offer insight into balancing fidelity and computational efficiency. Architectural optimizations could reduce resource usage while preserving accuracy, enhancing scalability. Integrating KANs with multimodal frameworks, e.g., facial expressions or gestures could enable advanced audio-visual synthesis, such as avatar generation, especially given the use of hypernetworks in visual domains (Przewięźlikowski et al., 2024; Kania et al., 2024). Developing interpretable models would improve transparency and trust, supporting real-world deployment and legal compliance. Lastly, enhancing compatibility with diverse audio encoding systems could boost reconstruction quality and expand KANs’ practical utility.
Appendix A Additional experiments
A.1. Extended ablation study
At the beginning of this section we conclude the ablation study (presented initially in Sect.  4.3). We show how KAN performance on 3 s audio samples is influenced by encoding length, layer sizes, spline order and  spline function grid size parameters. We consider the same hyperparameters as used for the 1.5 s experiments.
Table 12 shows the impact of different encoding lengths on 3 s utterances. The absolute performance is generally lower across all metrics in relation to 1.5 s samples, e.g. the best PESQ decreasing from 3.62 to 2.90 and the best SI-SNR decreasing from 25.97 to 17.17. The results are much less conclusive than in case of 1.5 s, where length 16 was clearly optimal across most metrics. However, the majority of metrics (MSE, PSNR, SI-SNR and PESQ, STOI) are the best for encoding lengths of 24 and 16, respectively.
Table 12. KAN performance on the 3 s samples, with respect to the encoding length.
Encoding #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@ A.1. Extended ablation study
2 23,016 3.6e-3 30.46 3.31 4.26 1.21 0.66 0.42
4 25,704 1.1e-3 35.41 2.03 10.31 1.55 0.82 0.29
8 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
10 33,768 6e-4 39.40 1.32 14.55 2.83 0.97 0.38
12 36,456 5e-4 39.66 1.37 14.79 2.78 0.98 0.37
16 41,832 4e-4 40.78 1.44 15.95 2.91 0.98 0.35
24 52,584 3e-4 41.93 1.60 17.18 2.84 0.98 0.34
Table 13 shows how sizes and depths of hidden layers influence results on 3 s samples. Similarly to 1.5 s (Table 8) the best performance is achieved by the largest network composed of [96, 48, 24] layers. The performance, however, does not improve with the increasing size as gradually as for 1.5 s samples.
Table 13. KAN results on 3 s samples in relation to different number and sizes of hidden layers.
Layer Sizes #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
[24, 12, 6] 10,500 7e-4 37.57 1.49 12.66 1.72 0.89 0.32
[24, 24, 12, 6] 18,564 6e-4 38.42 1.38 13.52 1.91 0.94 0.38
[48, 12, 6] 19,908 8e-4 38.01 1.33 13.03 2.16 0.94 0.37
[48, 12, 12, 6] 21,924 6e-4 38.31 1.34 13.40 2.18 0.95 0.39
[48, 16, 8] 23,408 7e-4 38.29 1.31 13.35 2.30 0.96 0.38
[48, 24, 12] 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
[96, 48, 24] 102,480 5e-4 40.96 1.18 16.12 3.52 0.99 0.16
In comparison to the 1.5 s case, where a spline order of size 3 is clearly the optimal hyperparameter setting, its advantage is less evident in Table 14. KAN with a spline order equal to 5 achieves higher PNSR and SI-SNR scores, without falling too much behind in PESQ and STOI metrics.
Table 14. KAN effectiveness in terms of diverse spline order sizes and 3 s long input
Spline #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
1 28,860 6e-4 38.50 1.47 13.61 2.29 0.95 0.33
2 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
3 33,300 8e-4 38.66 1.25 13.68 2.56 0.97 0.36
5 37,740 6e-4 39.09 1.27 14.19 2.45 0.97 0.34
7 42,180 6e-4 38.88 1.32 13.97 2.20 0.96 0.34
The results of the grid size parameter (displayed in Table 15) provide similar conclusions and tendencies as in the 1.5 s scenario (Table 10) — increasing the grid size leads to the improvement in all metrics but LSD.
Table 15. KAN effectiveness in terms of diverse grid size and 3 s long input
Grid size #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
1 11,100 2.9e-3 31.34 2.00 5.46 1.38 0.77 0.40
5 19,980 7e-4 38.11 1.47 13.21 1.97 0.92 0.24
10 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
12 35,520 5e-4 39.49 1.27 14.64 2.67 0.97 0.32
17 46,620 6e-4 39.73 1.32 14.81 3.07 0.97 0.27
Table 16. Effect of various learning rate values on 1.5 s and 3 s recordings.
Learning #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
Sample Length: 1.5 s
0.01 31,080 7e-4 40.56 1.37 19.47 3.28 0.99 0.19
0.007 31,080 7e-4 40.52 1.33 19.36 3.39 0.99 0.19
0.006 31,080 7e-4 40.02 1.31 18.92 3.35 0.99 0.18
0.005 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
0.001 31,080 1.1e-3 37.89 1.36 16.63 2.60 0.97 0.23
Sample Length: 3 s
0.01 31,080 6e-4 39.50 1.32 14.64 2.56 0.97 0.39
0.007 31,080 6e-4 39.23 1.28 14.34 2.57 0.97 0.37
0.006 31,080 6e-4 39.10 1.28 14.22 2.55 0.97 0.36
0.005 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
0.001 31,080 7e-4 37.76 1.47 12.80 1.93 0.93 0.30
At the end of this paragraph, we include ablation studies for the remaining hyperparameters. Table 16 shows the comparison of learning rates. A value of 0.001 yields the worst scores across all entries in the table, with the exception of the CDPAM metric. A learning rate of 0.01 achieves the best results in the PSNR and SI-SNR metrics, which are quantitative, while a learning rate near 0.006 produces the highest perceptual scores. Generally, KAN performs well and stably with higher learning rate setting. It can be considered its advantage over networks like SIREN, which are sensitive to the choice of the learning rate and their learning process can collapse when too high values are used.
Table 17 shows the comparison of performance of KAN depending on whether the enable scale spline flag is set to False or True. The former results in skipping the theoretically redundant weights. However, it leads to a drop in all metric scores except MSE, with perceptual metrics being affected the most.
Table 17. Effect of enabling the scale spline flag for 1.5 s and 3 s recordings.
Scale #Params MSE PSNR LSD SI-SNR PESQ STOI CDPAM
@@
Sample Length: 1.5 s
False 28,860 7e-4 39.41 1.53 18.23 2.74 0.97 0.28
True 31,080 7e-4 40.10 1.29 18.95 3.36 0.99 0.19
Sample Length: 3 s
False 28,860 7e-4 38.10 1.48 13.23 2.20 0.94 0.42
True 31,080 6e-4 38.93 1.27 14.03 2.50 0.97 0.36
A.2. INR
To show that KAN performs well across diverse types of audio, we train INRs on samples from the two aforementioned datasets - GTZAN (Tzanetakis and Cook, 2002) and UrbanSound8K (Salamon et al., 2014). We chose 100 samples for each dataset, 10 samples per genre in case of GTZAN and 10 samples per split in case of UrbanSound8K. The recordings were pre-processed in an analogous way to our main experiment - resampled to 22.05 kHz and cropped to 32,768 samples, resulting in audio of length 1.5 s. The results are presented in Tables 18 and 19. While FINER achieves the best PSNR and SI-SNR, it is the worst with respect to LSD. WIRE shows the best overall performance when all metrics are taken into account, followed by SIREN and KAN, both achieving similar results.
Table 18. Reconstruction metrics for GTZAN dataset. We report mean and standard deviation across 100 samples, 10 samples per genre.
INR (# params) MSE PSNR LSD SI-SNR